OpenStudy (anonymous):

(-2,-1) and (4,2) write an equation in slope-intercept form the line that passes through given points.

OpenStudy (anonymous):

m=3/6=1/2 y=1/2x+b 2=(1/2)(4)+b 2=2+b b=0 y=1/2x

OpenStudy (anonymous):

y=(1/2)x is the answer by the way